OpenStudy (anonymous):

What is the distance between point P(–2, 1, 3) and point Q(–1, 4, –2) to the nearest tenth?

OpenStudy (zzr0ck3r):

the different of P(a,b,c) and Q(d,e,f) is given by \[\sqrt{(a-d)^2+(b-e)^2+(c-f)^2}\]

OpenStudy (zzr0ck3r):

do you not understand?

OpenStudy (anonymous):

not really

OpenStudy (zzr0ck3r):

\[\sqrt{(-2-(-1))^2+(1-4)^2+(3-(-2))^2}=\sqrt{(-1^2)+(-3)^2+(5^2)}=\\\sqrt{1+9+25}=\sqrt{35} \approx 5.9\]

OpenStudy (zzr0ck3r):

ask a question if you don't understand